# Hypergrah-Enhanced Dual Convolutional Network for Bundle Recommendation

### v1

@@ Abstract
Bundle recommendations strive to offer users a set of items as a package named bundle, enhancing convenience and contributing to the seller’s revenue. While previous approaches have demonstrated notable performance, we argue that they may compromise the ternary relationship among users, items, and bundles. This compromise can result in information loss, ultimately impacting the overall model performance. To address this gap, we develop a unified model for bundle recommendation, termed hypergraph-enhanced dual convolutional neural network (HED). Our approach is characterized by two key aspects. Firstly, we construct a complete hypergraph to capture interaction dynamics among users, items, and bundles. Secondly, we incorporate U-B interaction information to enhance the information representation derived from users and bundle embedding vectors. Extensive experimental results on the Youshu and Netease datasets have demonstrated that HED surpasses state-of-the-art baselines, proving its effectiveness. In addition, various ablation studies and sensitivity analyses revealed the working mechanism and proved our effectiveness. Codes and datasets are available at https://github.com/AAI-Lab/HED
Index Terms:
Bundle Recommendation, Graph Convolutional Network, Hypergraph, Recommender System
I Introduction
@@
Recommender system is a type of system that automatically connects users with items, typically found on online platforms [1, 2, 3, 4, 5], assisting users in sorting and filtering items of interest from a vast selection [6]. In addition to recommending individual items to users, online service platforms such as Steam [^1] https://store.steampowered.com/ and Epic [^2] https://store.epicgames.com/, the digital distribution platforms for video games, and NetEase Cloud Music [^3] https://music.163.com/ and Spotify [^4] https://open.spotify.com/, the online music platforms, will also recommend bundles consisting of a number of items that have certain internal relationships. The prevalence of bundled items makes bundle recommendation an important task, and according to Deng et al.’s analysis of purchase statistics, more than 65.00% of game Love is Justice revenue in mobile games comes from these discounted bundles [7]. Therefore, bundle recommendation is an important part of the recommendation system. Bundle recommendation aims to improve the accuracy of the recommendation of bundles by mining the hidden messages of the interaction between users, items, and bundles.
@@
Compared to traditional recommendation, the difficulty of bundle recommendation is how to utilize item elements including item-user interaction and item-bundle interaction as auxiliary information for bundle recommendation. This brings two problems, how to introduce item elements with item-user interactions and item-bundle interactions as auxiliary information, and how to utilize auxiliary information to improve bundle recommendation performance. While there has been a lot of work in the area of bundle recommendations [8, 9, 10, 11, 12, 13, 14, 15], scholars’ understanding of constructing a triadic relationship of user-item-bundle has been biased. The left part of the figure 1 shows the understanding of the previous work[8, 9, 12, 13], specifically, they simplified the ternary relationship by arguing that bundle recommendations can be generalized along two paths from users to bundles and users to items to bundles. This understanding is based on the user’s perspective, but it tears down the user-item-bundle ternary relationship. This way of understanding lengthens the information transmission path and brings information loss. We believe that users, items, and bundles are a kind of mutual connection relationship, and there is intra-connection within users and bundles, and the right part of the figure 1 illustrates our understanding. However, current research lacks a user-item-bundle triad construction that matches our understanding, and this construction also increases the complexity of the model.
@@
Figure 1: Different understandings of the construction of the user-item-bundle triad: the right side is the understanding from previous work, and the left side is our understanding; where the blue squares represent users, the red squares represent items, and the green squares represent bundles; the grey arrows and the gold arrows represent the two paths from the understanding from previous work, and the black arrows and the bi-directional arrows represent the mutual connection; the blue curve and the green curve represent the intra connection of users and the intra connection of bundles, respectively.
@@
To this end, we design the hypergraph-enhanced dual convolutional neural network (HED) model to address the problem of how to introduce item elements and how to utilise them. The model comprises three parts: embedding initialization layer, dual convolutional layer, and prediction layer. Based on the work [14], we constructed the hypergraph based on the mutual connections between users, projects and bundles, as well as the internal connections between users and bundles. We termed the hypergraph created as a complete hypergraph. The embedding initialization layer generates embedding vectors for each user, item, and bundle. The dual convolutional layer is capable of extracting information from the complete hypergraph and merging the information from the U-B graph produced by U-B interaction to improve the embedding vector and prediction. The layer is suggested based on the ultimate embedding vectors of both users and bundles. Experiments on NetEase and Youshu show that our proposed method is ahead of the state-of-the-art methods.
@@
  • We construct a complete hypergraph to express the ternary relationship between users, items and bundles. The ternary relationship includes not only the mutual connections between users, items and bundles, but also the intra-connections within users and bundles.
@@
  • We conducted extensive experiments on two real-world datasets, and the results show that our proposed model outperforms the existing state-of-the-art baselines in terms of Recall and NDCG evaluation metrics in all three scenarios: Top 20, Top 40, and Top 80.
II Related Works
@@
As an increasing number of scholars focus on bundle recommendation, the field has experienced rapid growth, giving rise to numerous advanced works. A portion of the work focuses on how to decouple the complex relationship between users, items, and bundles. For example, Chen et al. [16] aggregated item embeddings within a bundle using an attention mechanism to obtain a representation of the bundle, and jointly modeled user-bundle interactions and user-item interactions with a multitasking approach. Tan et al. [9] used a hierarchical approach to disentangle user preferences for items and bundle preferences. In contrast, Avny et al. [11] improved the Transformer structure to represent the data of users, items, and bundles, and achieved modeling of users’ preferences for each item in the bundle and for the whole bundle through a self-attention mechanism.
@@
Influenced by graph neural networks in the field of recommender systems, many works have chosen to represent the relationships among users, items, and bundles with graph structures, transforming bundle recommendation into a link prediction problem on graphs. For example, Chang et al. [8] used two heterogeneous graphs to represent the three kinds of interactions among users, items, and bundles, and employed graph convolutional propagation to learn the user and bundle representations. Wang et al. [10] augmented the representations characterizing users and bundles with neighbor interactions. Ma et al. [12] represented the relationship between the three with bundle views and project views, and chose to use comparative learning to obtain user and bundle representation. Zhao et al. [13] learned the user and bundle representations by comparing the user intention in the global view and the user intention in the local view with contrast learning. Yu et al. [14] constructed the three kinds of interactions into a unified hypergraph and used a graph convolutional neural network to learn the user and bundle representations. Wei et al. [15] used the graph transformer to model and learn bundle strategies, and combined user preferences for items to obtain representations of users and bundles. These works aim to capture the three interactions involving users, items, and bundles using graph-based methodologies. In particular, the study by Yu et al. [14] utilized hypergraphs to represent these interactions. However, it has been observed that this representation is not comprehensive and does not sufficiently emphasize the role of items in bundle recommendations. To address this limitation, we have undertaken further refinement and construction of complete hypergraphs, building upon the foundation laid out in [14]. Our objective is to create a more detailed and unified structure that effectively encapsulates the interactions between users, items, and bundles, as well as the relationships within these three components.
III Problem Formulation
@@
Consider a dataset with three main entities: bundles represented by the set $S_{B}=\{b_{i}\mid i\in 1,2,\dots,B\}$, users represented by the set $S_{U}=\{u_{j}\mid j\in 1,2,…,U\}$, and items represented by the set $S_{I}=\{i_{k}\mid k\in 1,2,\dots,I\}$. The interactions within this dataset are captured through three matrices:
  • User’s purchase interactions with bundles are represented by the matrix $A_{ub}$, where each entry $A_{ub}(i,j)$ signifies the user $u_{j}$’s interaction with bundle $b_{i}$.
  • User’s purchase interactions with individual items are captured by the matrix $A_{ui}$, with each entry $A_{ui}(j,k)$ indicating whether user $u_{j}$ has purchased item $i_{k}$.
  • The relationships between bundles and items are represented by the matrix $A_{bi}$, where each entry $A_{bi}(i,k)$ indicates whether bundle $b_{i}$ includes item $i_{k}$ as a subordinate component.
@@
Given the assumptions, the task of bundle recommendation can be formulated as predicting the likelihood of purchasing bundles based on purchase interactions and subordinate interactions. Specifically, the model takes as inputs the matrices representing purchase interactions ($A_{ub}$ and $A_{ui}$), and subordinate interactions ($A_{bi}$). The output of the model is a set of personalized score values that indicate a user’s preference for purchasing bundles. The personalized scoring function is expressed as follows:
@@
IV Methodology
The formulated problem is addressed through the framework of the hypergraph-enhanced dual convolutional neural network (HED for short). The framework’s structure is depicted in Figure 2, highlighting key components such as the Embedding Initialization Layer, the Dual Convolutional Layer, and the Prediction Layer.
To elucidate the construction of the hypergraph in this paper, refer to Section IV-A. This section outlines the methodology for building the complete hypergraph that is central to the proposed model. Additionally, Section IV-B provides insights into the proposed graph neural network model.
Figure 2: The framework graph of HED: 1) Embedding initialization layer, which corresponds to subsection IV-B1 of the article; 2) Dual Convolutional Layer, which corresponds to subsection IV-B2 of the article; and 3) Prediction Layer, which corresponds to subsection IV-B3 of the article, blue squares represent the user, red squares represent the item, and green squares represent the bundle; rectangles of different colors represent the embedding vectors of different elements, solid arrows of different colors represent the data transfer of different elements, and solid arrows of different colors represent the data updating of different elements, blue represents user, red represents item, and green represents bundle, the construction method of the hypergraph $H$ corresponds to subsection IV-A of the article, the U-B graph is denoted as $h_{hb}$ in the article, the dashed box and ellipsis in the middle of the two images represent the omitted elements and interactions, the ellipsis in the embedding section represents the embedding corresponding to the omitted elements, and the dashed box in the Prediction Layer section represents the generated recommendation sequence..
IV-A Complete Hypergraph Structure
To address this limitation, the work [14] introduced the concept of hypergraphs to bundle recommendation, offering a more expressive framework capable of connecting any number of vertices with a single hyperedge. The proposed unified hypergraph construction method in [14] lays a foundation for capturing the interplay between users, items, and bundles. However, it has notable gaps. Specifically, the unified hypergraph construction method overlooks the inherent interactions of each user and each bundle with itself. Additionally, it underplays crucial aspects such as the purchase interactions of users with items and the affiliation interactions of bundles with items.
@@ IV-A Complete Hypergraph Structure
Building upon the foundation laid by [14], this paper seeks to enhance hypergraph construction by introducing complete hypergraphs. Figure 3 provides a visual representation of these complete hypergraphs. In contrast to the unified hypergraph, our approach meticulously incorporates self-interactions of users and bundles, while also accounting for the intricate dynamics of user-item purchase interactions and bundle-item affiliation interactions.
Figure 3: Schematic representation of the hypergraph synthesis method: blue squares indicate users, red squares indicate items and green squares indicate bundles. Black arrows indicate computation and gold arrows indicate synthesis. By interacting $A_{ub}$, $A_{ui}$ and $A_{bb}$, U-B graph $h_{ub}$, U-I graph $h_{ui}$ and B-I graph $h_{bi}$ can be generated. The U-U graph (denoted as $h_{uu}$) and the B-B graph (denoted as $h_{bb}$) are calculated based on the interaction $h_{ub}$. In addition, $h_{ii}$ is the zero matrix and the ellipses represent omitted users, items, bundles, and corresponding interactions. The hypergraph $H$ is formed by the combination of these interaction graphs, denoted by the formula 2.
In the figure, the B-U graph also referred to as the B-U interaction matrix ($h_{ub}$), is derived from the interaction matrix $A_{ub}$. To clarify, the nodes in $h_{bu}$ correspond to the interactions between users and bundles. If $(i,j)$ is present in $A_{ub}$, it means that user $u_{i}$ purchased bundle $b_{j}$, so $h_{ub}(i,j)=1$, and vice versa $h_{ub}(i,j)=0$. Essentially, both $A_{ub}$ and $h_{ub}$ are mathematical representations of the relationship between a user and a bundle that generates a purchase, where $A_{ub}$ is represented by a pair of numbers representing the user number and the bundle number, and $h_{ub}$ is represented by the graph structure $h_{ub}(i,j)=1$. Similarly, the interaction $A_{ui}$ and $A_{bi}$ can be obtained as the U-I graph $h_{ui}$ and B-I graph $h_{bi}$. The matrix $h_{bu}$ is the transpose of the $h_{ub}$ matrix. Similarly, $h_{iu}$ and $h_{ib}$ are the transpose matrices of $h_{ui}$ and $h_{bi}$, respectively. The matrices $h_{uu}$ and $h_{bb}$ signify user-user and bundle-bundle interactions, respectively. These can be derived from the $h_{ub}$ interaction matrix. Specifically, for users A and B, if the number of bundles with the same interaction is greater than $n$, then users A and B are considered to have an interaction. This corresponds to $h_{uu}(a,b)$ and $h_{uu}(b,a)$ in the $h_{uu}$ matrix being set to 1, and vice versa. The parameter $n$ is experimentally tested under different datasets and is consistent with  [14]. Furthermore, it is important to note that each user is considered to interact with themselves. Consequently, we set the elements on the main diagonal of $h_{uu}$ to 1. The same principle applies to $h_{bb}$. Regarding $h_{bb}$, despite experimenting with construction methods similar to those used for $h_{uu}$ and attempting to set it as a unit matrix, the optimal experimental results were achieved when constructing $h_{bb}$ as a zero matrix. The complete hypergraph construction method can be represented by the formula 2, while the formula 3 is the construction method of the unified hypergraph [14]. In contrast to the unified hypergraph $H_{Unify}$, our construction of the full hypergraph $H$ emphasizes the interaction of items with bundles and users and can fully express the understanding of the right-hand side of the graph 1 concerning bundle recommendations.
@@
$H_{Unify}=\begin{bmatrix}h_{ui}&h_{uu}&h_{ub}\\ h_{bi}&h_{bu}&h_{bb}\\ \end{bmatrix}$ (3)
IV-B Network Structure
IV-B1 Embedding Initialization Layer
The focus of this paper lies in the domain of implicit feedback, where the available information is confined to the easily accessible history of interactions. Given this constraint, our model relies on the random initialization of embedding vectors for users, items, and bundles. These initialized embeddings constitute a crucial component of the trainable parameters within the model.
Building on the construction method outlined for the complete hypergraph, we arrange the embedding vectors to serve as inputs for the subsequent modules of the model. The stacking of these embedding vectors is carried out in accordance with the construction principles of the complete hypergraph.
Assuming that the embedding size of the model is set to $d$, the initialization generates an embedding of $E_{u}^{(1)}\in R^{U\times d}$ for users, $E_{i}^{(1)}\in R^{I\times d}$ for items, and $E_{b}^{(1)}\in R^{B\times d}$ for bundles, after the stacked embedding vector $E^{(1)}\in R^{(U+I+B)\times d}$, the formula for the stacking method is expressed as follows.

### v2

@@ Abstract
Bundle recommendations strive to offer users a set of items as a package named bundle, enhancing convenience and contributing to the distribution system’s revenue. Although existing bundle recommendation methodologies have demonstrated the ability to effectively capture user, item and bundle interactions, we posit that these approaches tend to compartmentalise the complex ternary relationship between users, items and bundles into two distinct paths, rather than considering it in its entirety. This compromise can result in information loss, ultimately impacting the overall model performance.
@@
To address this gap, we develop a unified model for bundle recommendation, termed Hypergraph-Enhanced Dual convolutional neural network (HED). Three key aspects characterize our approach. First, we construct a complete hypergraph to capture the interactions dynamics between users, items, and bundles. Second, to enhance the quality of user and bundle embeddings and better utilize the complete hypergraph constructed, we generate embeddings for each item in addition to those for each user and bundle. Third, influenced by the coarse and fine ranking of recommender systems, we design a dual convolution module to achieve more accurate recommendations. Extensive experimental results on the Youshu and Netease datasets have demonstrated that HED surpasses state-of-the-art baselines, proving its effectiveness. Various ablation studies and sensitivity analyses revealed the working mechanism and proved our effectiveness. In addition, we discuss hypergraph construction and implementation efficiency. Codes and datasets are available at https://github.com/AAI-Lab/HED
keywords:
[…]
1 Introduction
A recommender system is a crucial component of a Internet distribution system that rapidly searches for items, such as books, music, videos, and games, that may be of interest to users from a vast collection of items [21, 23, 22, 8, 52]. The fundamental principle of recommender systems is to recommend items that are relevant to users, employing feature engineering techniques that utilise user preferences, item characteristics, and user interactions (purchases or clicks) [20].
@@ 1 Introduction
In addition to recommending individual items to users, online service platforms such as Steam [^1] https://store.steampowered.com/, Epic [^2] https://store.epicgames.com/, the digital distribution platforms for video games, and NetEase Cloud Music [^3] https://music.163.com/, Spotify [^4] https://open.spotify.com/, the online music platforms, will also recommend bundles consisting of a number of items that have certain internal relationships. The prevalence of bundles has made bundle recommendation an important task, according to Deng et al.’s analysis of purchase statistics, more than 65.00% of game Love is Justice revenue in mobile games comes from these discounted bundles [10]. Bundle recommendation aims to improve the accuracy of the recommendation of bundles by mining the hidden messages of the interactions between users, items, and bundles.
@@
Compared with traditional recommendation, the difficulty of bundle recommendation is how to introduce and utilize item information to improve the performance of bundle recommendation. This brings up two questions, how to structure the user-item-bundle triadic relationship, and how to leverage the triadic relationship and focus on users and bundles for recommendations. While there has been a lot of work in the area of bundle recommendations [7, 32, 39, 2, 28, 51, 49, 42], scholars understanding of the triadic relationship of user-item-bundle has been biased. The left part of the figure 1 shows the understanding of the previous work[7, 32, 28, 51]. They simplified the ternary relationship of bundles by arguing that bundle recommendations can be generalized along two paths: path 1 (users-bundles) and path 2 (users-items-bundles). This understanding artificially dismantles the user-item-bundle triadic relationship from the user’s point of view, lengthens the information transfer path and introduces information loss. We believe that users, items and bundles are in an interconnected relationship (both in terms of purchase and affiliation), and that there are internal connections within users or bundles, as illustrated by the right part of the figure 1. However, the current research does not conform to our understanding of the user-item-bundle ternary structure, nor does it have a bundle recommendation model based on this design.
@@
Figure 1: Different understandings of the construction of the user-item-bundle triadic relationship: the left side is the understanding from previous work, and the right side is our understanding. The blue, red, and green squares represent users, items, and bundles, respectively. The grey and gold arrows represent the two types of paths from the understanding from previous work, while the black and bi-directional arrows represent the mutual connection from our understanding. The blue and green curves represent the intra connection of users and the intra connection of bundles, respectively.
@@
To this end, we design the Hypergraph-Enhanced Dual convolutional neural network (HED) model. The model comprises three parts: Embedding initialization layer, Dual convolutional layer, and Prediction layer. Based on the work [49], we constructed hypergraphs called complete hypergraphs based on the mutual connections between users, items, and bundles as well as the internal connections between users or bundles. The Embedding initialization layer generates embeddings for each user, item, and bundle, with the aim of improving the quality of user and bundle embeddings through item embeddings. The Dual convolutional layer is composed of two parts. The first part convolves the hypergraph to obtain information about the user, item, and bundle from the complete hypergraph. The second part convolves the U-B graph, which is obtained from the U-B interactions, to obtain information about the user and the bundle. The final embedding of the user and the bundle is obtained by combining the information from both parts. The Prediction layer is suggested based on the ultimate embedding of both users and bundles. Experiments on NetEase and Youshu show that our proposed method is ahead of the state-of-the-art methods.
@@
  • We construct a complete hypergraph to express the ternary relationship between users, items, and bundles. The ternary relationship includes not only the mutual connections between users, items, and bundles but also the intra connections within users or bundles.
@@
  • We conducted extensive experiments on two real-world datasets, and the results show that our proposed model outperforms the existing state-of-the-art baselines in terms of Recall and NDCG evaluation metrics in all three scenarios: Top 20, Top 40, and Top 80. Furthermore, the increase in NDCG metrics and the metrics of Top 20 demonstrates that the HED model enhances the precision of recommendations.
2 Related Works
This section presents a review of related work in three areas: (1) Graph-Based Recommender System, (2) Hypergraph-Based Recommender System, and (3) Bundle Recommendation.
2.1 Graph-Based Recommender System
A recommender system is a system that learns from a user’s historical interactions with items and subsequently recommends items to the user that may be of interest to them. The user and the item can be regarded as nodes of a graph, with the interaction information between them forming the edges of the graph. This interaction information is thus well-suited to expression in the form of a graph. The development of graph neural networks has led to the clarification of several key paradigms in the field of graph analysis, including link prediction, graph classification, graph generation, and graph node classification. Additionally, the emergence of classical graph neural network algorithms, such as Spectral CNN[5], GCN[19], GraphSAGE[16], and GAT[33], has facilitated the development of graph neural network-based recommendation models.
In the field of personalised recommendation, He et al. innovatively constructed the interaction information of users and items as a bipartite graph, and designed the Neural Graph Collaborative Filtering framework, which learns the higher-order information of the bipartite graph and updates it [37]. The embedding of users and items through GCN. NGCF not only demonstrates outstanding performance, but also sets a precedent for GNN-based recommender systems. He et al. demonstrated that feature transformation and nonlinear activation can negatively affect GCNs through ablation analysis, which led to the development of the far-reaching LightGCN model as the research progressed [17]. Liu et al. decomposed the embedding into multiple segments related to semantics, thereby enhancing the performance and interpretability of the model [25]. Wu et al. enhanced accuracy and robustness through self-supervised learning based on LightGCN, thereby alleviating the long-tail problem [43]. Xia et al. proposed an incremental graph convolutional network (GCN) as a means of addressing the cold-start problem inherent to GCN when applied to new users and new items [44]. Lin et al. enhanced the efficacy of recommendation algorithms by integrating contrast learning, which categorizes nodes and their structural neighbors as positive contrast pairs [24].
The prevalence of graph neural networks is not only limited to the field of personalised recommendation, but has also flourished in the fields of hybrid recommendation  [15], session recommendation  [31], point-of-interest (POIs) recommendation  [41], group recommendation  [1], bundle recommendation [28], and other segmented recommender system domains have also flourished. Gatzioura et al. developed a graph-based model that incorporates content-based and collaborative filtering-based models for mining music descriptions with user preferences for music recommendation [15]. Sheng et al. constructed interactions and sessions as Weighted Global Item Graph and Local Session Graph, respectively, and designed an Enhanced Graph Neural Network to mine the information of the two graphs for session recommendation [31]. Wang et al. put forward the Adaptive Graph Representation-enhanced Attention Network (AGRAN) framework with a view to learning spatio-temporal information of POI and the user’s preferences for POI recommendation, using a combination of techniques including graph structure learning and the construction of an adaptive POI graph matrix [41]. Abolghasemi et al. evaluated the similarity of user graphs using a graph convolutional neural network and designed the GcPp technique to cluster highly similar users into clusters, which in turn enables group recommendation [1]. Furthermore, information external to users and items can be represented as a knowledge graph, which in turn facilitates the generation of more accurate recommendations based on historical interactions [13].
2.2 Hypergraph-Based Recommender System
As the study progressed, the use of simple graphs was no longer sufficient to convey the required information. Hypergraph has attracted significant attention due to its capacity to accommodate diverse data types and establish intricate connections, offering a promising avenue for addressing the challenges posed by such vast and heterogeneous datasets [4]. The proposal of hypergraph neural networks has led to the emergence of a growing body of research on hypergraph-based approaches in the field of recommender systems [12, 14]. This includes studies on personalized recommendation [35, 18, 45, 40, 27], POI recommendation [47], session-based recommendation [46, 11], group recommendation [50, 36], and social recommendation [48].
In a personalized recommendation task, Wang et al. constructed a series of hypergraphs based on item correlations for different time periods and employed a Hypergraph Convolution Network combined with self-attention to learn hypergraph information for recommendation [35]. The majority of works [18, 45, 40, 27] construct hypergraphs based on users and items. Ji et al. designed a two-channel collaborative filtering framework to learn hypergraph information [18]. Xia et al. constructed hypergraphs based on users and items and made recommendations through contrastive learning [45]. Wang et al. constructed hypergraphs based on users and items and made recommendations through attention and dynamics clustering [40]. Ma et al. constructed a hypergraph based on users and items and learned hypergraph information through cross-view hypergraph contrastive learning for recommendation [27].
In the session-based recommendation task, scholars construct hypergraphs based on sessions and items, Xia et al. learn hypergraph information through a two-channel hypergraph convolutional network and train the model through a self-supervised task [46], and Ding et al. learn two item representations through Hypergraph Convolutional Networks (HGCN) and Gate Recursive Units (GRU) with an attentional mechanism for recommendation [11].
In the social recommendation task, Yu et al. constructed a hypergraph based on social network and user-item interaction graph, and used self-supervised learning to train the hypergraph convolutional network [48]. In the POI recommendation task, Yan et al. constructed a trajectory-based hypergraph and proposed a spatio-temporal hypergraph convolutional network that fuses complex higher-order information and global collaborative relationships among trajectories for recommendation [47].
2.3 Bundle Recommendation
@@
As an increasing number of scholars focus on bundle recommendation, the field has experienced rapid growth, giving rise to numerous advanced works. A portion of the work focuses on how to decouple the complex relationship between users, items, and bundles. For example, Rendle et al. treat bundles as atomic, use matrix decomposition to learn preferences, and make recommendations via Markov chains [30]. Similarly, Cao et al. employed a factor decomposition model to capture user preferences for items and bundles and to make recommendations based on co-occurring information about items and bundles [6]. In addition, the method they developed addresses the cold-start problem for new items. Chen et al. [9] aggregated item embeddings within a bundle using an attention mechanism to obtain a representation of the bundle, and jointly modeled user-bundle interactions and user-item interactions with a multitasking approach. Tan et al. [32] used a hierarchical approach to disentangle user preferences for items and bundle preferences. In contrast, Avny et al. [2] improved the Transformer structure to represent the data of users, items, and bundles, and achieved modeling of users’ preferences for each item in the bundle and for the whole bundle through a self-attention mechanism.
Figure 2: The framework graph of HED: 1) Embedding initialization layer, which corresponds to subsection 4.2.1 of the article; 2) Dual Convolutional Layer, which corresponds to subsection 4.2.2 of the article; and 3) Prediction Layer, which corresponds to subsection 4.2.3 of the article. The blue, red, and green squares represent users, items, and bundles, respectively. Rectangles of different colors represent the embedding of different elements. Data transmission and updating are shown by solid and dashed lines. The hypergraph H is the complete hypergraph constructed by subsection 4.1. The dashed boxes in the prediction layer section indicate the generated recommendation sequences.
@@
Influenced by graph neural networks in the field of recommender systems, many works have chosen to represent the relationships among users, items, and bundles with graph structures, transforming bundle recommendation into a link prediction problem on graphs. For example, Chang et al. [7] used two heterogeneous graphs to represent the three kinds of interactions among users, items, and bundles, and employed graph convolutional propagation to learn the user and bundle representations. Wang et al. [39] augmented the representations characterizing users and bundles with neighbor interactions. Ma et al. [28] represented the relationship between the three with bundle views and item views, and chose to use comparative learning to obtain user and bundle representation. Zhao et al. [51] learned the user and bundle representations by comparing the user intention in the global view and the user intention in the local view with contrast learning. Yu et al. [49] constructed the three kinds of interactions into a unified hypergraph and used a graph convolutional neural network to learn the user and bundle representations. Wei et al. [42] used the graph transformer to model and learn bundle strategies, and combined user preferences for items to obtain representations of users and bundles. These works aim to capture the three interactions involving users, items, and bundles using graph-based methodologies. In particular, the study by Yu et al. [49] utilized hypergraphs to represent these interactions. However, this approach is not complete enough. Although items are introduced, the hypergraph underutilizes the interaction information for the included items resulting in an incompletely constructed hypergraph, and there is no generation of embedding for items to assist users and bundles in learning the hypergraph structure, as well as a lack of highlighting users interacting with bundles after learning the hypergraph. These factors not only led to poor model performance with the larger and more complex NetEase dataset but also led to the appearance of bundles that users wanted to interact with in a backward position.
3 Problem Formulation
@@
Consider a dataset with three main entities: bundles represented by the set $S_{B}=\{b_{i}\mid i\in 1,2,…,B\}$, users represented by the set $S_{U}=\{u_{j}\mid j\in 1,2,…,U\}$, and items represented by the set $S_{I}=\{i_{k}\mid k\in 1,2,…,I\}$. The interactions within this dataset are captured through three matrices: User’s purchase interactions with bundles are represented by the matrix $A_{ub}$, where each entry $A_{ub}(i,j)$ signifies the user $u_{j}$’s interactions with bundle $b_{i}$. The same reasoning can be used to obtain the purchase interactions of users with individual items $A_{ui}$ and the affiliation interactions of bundles and items $A_{bi}$.
@@
Given the assumptions, the task of bundle recommendation can be formulated as predicting the likelihood of purchasing bundles based on purchase interactions and affiliation interactions. Specifically, the model takes as inputs the matrices representing purchase interactions ($A_{ub}$ and $A_{ui}$), and affiliation interactions ($A_{bi}$). The output of the model is a set of personalized score values that indicate a user’s preference for purchasing bundles. The personalized scoring function is expressed as follows:
Figure 3: Schematic diagram illustrating the structure of the hypergraph: blue squares indicate users, red squares indicate items and green squares indicate bundles. Black arrows indicate computation and gold arrows indicate synthesis. By interacting $A_{ub}$, $A_{ui}$ and $A_{bb}$, U-B graph $h_{ub}$, U-I graph $h_{ui}$ and B-I graph $h_{bi}$ can be generated. The U-U graph (denoted as $h_{uu}$) and the B-B graph (denoted as $h_{bb}$) are calculated based on the interactions $h_{ub}$. In addition, $h_{ii}$ is the zero matrix and the ellipses represent omitted users, items, bundles, and corresponding interactions. The hypergraph $H$ is formed by the combination of these interaction graphs, denoted by the Equation (2).
4 Methodology
The formulated problem is addressed through the framework of the HED. The framework’s structure is depicted in Figure 2, highlighting key components such as the Embedding Initialization Layer, the Dual Convolutional Layer, and the Prediction Layer.
The construction of the hypergraph is explained in section 4.1. This section outlines the methodology for building the complete hypergraph that is central to the proposed model. Additionally, Section 4.2 provides insights into the proposed graph neural network model.
4.1 Complete Hypergraph Structure
@@
To address this limitation, the work [49] proposes a unified hypergraph construction methodology that introduces the concept of hypergraphs to bundle recommendations, laying the groundwork for capturing interactions between users, items, and bundles. However, it has notable gaps. Specifically, the unified hypergraph construction method overlooks the inherent interactions of each user and each bundle with itself. Additionally, it underplays crucial aspects such as the purchase interactions of users with items and the affiliation interactions of bundles with items.Building upon the foundation , this paper seeks to enhance hypergraph construction by introducing complete hypergraphs. Figure 3 provides a visual representation of these complete hypergraphs. In contrast to the unified hypergraph, our approach meticulously incorporates self-interactions of users or bundles, while also accounting for the intricate dynamics of user-item purchase interactions and bundle-item affiliation interactions.
@@
In the figure, the B-U graph also referred to as the B-U interaction matrix ($h_{ub}$), is derived from the interaction matrix $A_{ub}$. To clarify, the nodes in $h_{bu}$ correspond to the interactions between users or bundles. If $(i,j)$ is present in $A_{ub}$, it means that user $u_{i}$ purchased bundle $b_{j}$, so $h_{ub}(i,j)=1$, and vice versa $h_{ub}(i,j)=0$. The matrices $h_{uu}$ and $h_{bb}$ signify user-user and bundle-bundle interactions, respectively. These can be derived from the $h_{ub}$ interaction matrix. Specifically, for users A and B, if the number of bundles with the same mutual interaction is greater than $n$, then users A and B are considered to have an intra interaction. This corresponds to $h_{uu}(a,b)$ and $h_{uu}(b,a)$ in the $h_{uu}$ matrix being set to 1, and vice versa. The parameter $n$ is experimentally tested under different datasets and is consistent with  [49]. Furthermore, it is important to note that each user is considered to interact with themselves. Consequently, we set the elements on the main diagonal of $h_{uu}$ to 1. The same principle applies to $h_{bb}$. Regarding $h_{ii}$, despite experimenting with construction methods similar to those used for $h_{uu}$ and attempting to set it as a unit matrix, the optimal experimental results were achieved when constructing $h_{ii}$ as a zero matrix. The specific analysis can be seen in 5.5. The complete hypergraph construction method can be represented by the Equation (2), our construction of the full hypergraph $H$ emphasizes the interactions of items with bundles and users and can fully express the understanding of the right-hand side of the graph 1 concerning bundle recommendations.
4.2 Network Structure
4.2.1 Embedding initialization layer
The focus of this paper lies in the domain of implicit feedback, where the available information is confined to the easily accessible history of interactions. Given this constraint, our model relies on the random initialization of embedding for users, items, and bundles. These initialized embeddings constitute a crucial component of the trainable parameters within the model.
Building on the construction method outlined for the complete hypergraph, we arrange the embedding to serve as inputs for the subsequent modules of the model. The stacking of these embedding is carried out in accordance with the construction principles of the complete hypergraph.
Assuming that the embedding size of the model is set to $d$, the initialization generates an embedding of $E_{u}^{(1)}\in R^{U\times d}$ for users, $E_{i}^{(1)}\in R^{I\times d}$ for items, and $E_{b}^{(1)}\in R^{B\times d}$ for bundles, after the stacked embedding vector $E^{(1)}\in R^{(U+I+B)\times d}$, the equation for the stacking method is expressed as follows.